OpenStudy (anonymous):

Is there anybody on here who is taking Algebra 2 Segment 1 on FLVS or has taken it already?

Nnesha (nnesha):

FLVS = flying vs. swimming ? post your question please

OpenStudy (anonymous):

lol @Nnesha

OpenStudy (anonymous):

For this adventure, you and world renowned Professor Sherlock McMerlock are traveling to the Lost Island of Laplaya. When the boat arrives on the island shore, you and Professor McMerlock disembark on your adventure. You trudge through the jungles and arrive at three impressively large doors. About eye level on each door is an intricately carved keyhole. Directions are scratched into the wood above each keyhole. To open these doors, you must match the number and type of solutions for the following two functions in standard form. f(x) = x2 + 6x – 16 g(x) = x2 +6x + 1 Match the following descriptions of the solutions to each of the functions above. Hint: they each have their own match. Two real irrationals solutions Two real rationals solutions

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

Im stuck on that assignment as well but i did solve the first equation.

OpenStudy (anonymous):

i used the functions f(x)=b^2-4ac and f(x)=-b+or- sqrt(b^2-4ac)) OVER 2a

OpenStudy (anonymous):

@Fallenangell16 I'm stuck on this assignment and I sicked around on FLVS so I'm 60% done but I still need help on this assignment

OpenStudy (anonymous):

im way behind on FLVS because i got pulled out of public school and had to restart the whole year over again so im just now on this assingment. I dont know if im doing it correctly but i could give you what i did to solve the first one.

OpenStudy (anonymous):

I think I might know how to do it @Fallenangell16

OpenStudy (anonymous):

okayy